# Streaming Technologies and Serialization Protocols: Empirical Performance Analysis

### v1

@@ Abstract
Efficiently streaming high-volume data is essential for real-time data analytics, visualization, and AI and machine learning model training. Various streaming technologies and serialization protocols have been developed to meet different streaming needs. Together, they perform differently across various tasks and datasets. Therefore, when developing a streaming system, it can be challenging to make an informed decision on the suitable combination, as we encountered when implementing streaming for the UKAEA’s MAST data or SKA’s radio astronomy data. This study addresses this gap by proposing an empirical study of widely used data streaming technologies and serialization protocols. We introduce an extensible and open-source software framework to benchmark their efficiency across various performance metrics. Our findings reveal significant performance differences and trade-offs between these technologies. These insights can help in choosing suitable streaming and serialization solutions for contemporary data challenges. We aim to provide the scientific community and industry professionals with the knowledge to optimize data streaming for better data utilization and real-time analysis.
@@ I Introduction
With the exponential increase in data generation from large scientific experiments and the concurrent rise of data-intensive machine learning algorithms within scientific computing [1], traditional methods of data transfer are becoming inadequate. The trend necessitates efficient data streaming methods that allow end-users to access subsets of data remotely. Additionally, the drive for FAIR and open data [2, 3] mandates that such data are, ultimately, publicly accessible to end users over a wide-area network connection.
The Mega-Ampere Spherical Tokamak (MAST) [4] was a spherical tokamak in operation at the UK Atomic Energy Authority (UKAEA), Culham Centre for Fusion Energy (CCFE) from 1999 to 2013 and its upgraded successor was the MAST-U [5] began operation in 2020. These facilities generate gigabytes of data per experimental shot, accumulating substantial data daily.
The lack of public accessibility to the historical archive of data produced by the MAST experiment has limited collaborative opportunities with international and industry partners. The pressing need to facilitate real-time data analysis [6] and leverage recent advancements in machine learning [7] further emphasizes the necessity for efficient data streaming technologies. These technologies must not only handle the sheer volume of data but also integrate seamlessly with analytical tools.
In this paper, we extend the work conducted in [8] for the SKA’s radio astronomy data streaming and visualization. We explore an array of streaming technologies available. We consider the combination of two major choices of technology when implementing a streaming service: (a) the choice of a streaming system, which performs the necessary communication between two endpoints, and (b) the choice of encoding used to convert the data into transmittable formats. Our contributions are as follows:
  • We provide a comprehensive review of 11 streaming technologies and 11 encoding methods, categorized by their underlying principles and operational frameworks.
@@
  • By testing 132 combinations, we offer a detailed comparative analysis of their performance across six different data types.
@@
Section II briefly reviews the related work in this area. Section III provides an overview of the different serialization protocols and data streaming technologies reviewed in this study. Section IV outlines our experimental methodology, including the performance metrics considered, implementation details of our benchmark framework and the choice of datasets used for evaluation. Section V discusses the results of our experiments across all performance metrics and datasets. Finally, sections VI and VII reflect on the results of our study and draw recommendations of technology choices.
@@ II Related Work
Churchill et al. [6] explored using ADIOS2 [14] for transferring large amounts of Tokamak diagnostic data from the K-STAR facility in Korea to the NERSC and PPPL facilities in the USA for near-real-time data analysis.
We differentiate our study from these related works by evaluating 1) A wide variety of different streaming technologies, both message broker-based and RPC-based. 2) considering a large number of data serialization formats, including text, binary, and protocol-based formats. 3) We evaluate the combination of these technologies, developing an extensible framework for measuring and comparing serialization and streaming technologies. 4) Evaluating the performance over 10 different metrics. We comprehensively evaluate 10 different streaming technologies with 12 different serialization methods over 8 different datasets.
@@ III-A Serialization Protocols
TABLE I: A comparison of various serialization protocols. Type: describes how the method serializes data, whether in text or binary format, or relying on a common protocol. Human readable: indicates whether the serialization scheme is legible to a human reader. Defined schema: specifies whether producer and consumer share a common knowledge of the data format prior to transmission. Code generated schema: states whether the serialization requires code to be generated from a predefined protocol.
Protocol Type Binary Human Readable Defined Schema Code Generated Schema Based On
@@
BSON [18] Binary JSON
UBSON [19] Binary JSON
CBOR [20] Binary MessagePack
MessagePack [21] Binary JSON
@@
Capn’Proto [25] Protocol
@@ III-A1 Text Formats
JavaScript Object Notation (JSON) [16] is another human-readable data interchange format that represents data as a collection of nested key-value pairs. JSON is commonly used for data exchange protocol in RESTful APIs. Due to the smaller payload size, it is often seen as a lower overhead alternative to XML for data interchange.
@@ III-A3 Protocol Formats
Thrift [24] is another binary data format developed by Apache Software Foundation or Apache, similar in many respects to ProtoBuf. In Thrift, data structures are also defined in a separate file, and these definitions are used to generate corresponding appropriate data structures in various supported languages. Before transmission, data is serialized into a binary format. Thrift is also designed for RPC communication and includes methods for defining services that use Thrift data structures. However, Thrift has a smaller number of supported data types compared to ProtoBuf.
@@
We also considered the PSON format [28] and Zerializer [29]. PSON is a binary serialization format with a current implementation limited to C++ and lacks Python bindings, which restricts its applicability for our study. Zerializer, on the other hand, necessitates a specific hardware setup for implementation, placing it outside the scope of our study due to practical constraints. Consequently, while these formats might offer potential advantages, their limitations in terms of language support and hardware requirements precluded their inclusion in our experimental evaluation.
A summary of serialization protocols can be found in Table I. The text-based formats represent data during a text-based markup. While human-readable, text-based formats suffer from larger payload and serialization costs due to the overhead of the markup describing the data. In contrast, binary formats serialize the data to bytes before transmission. These formats are not human-readable, but achieve a better payload size with lower serialization costs. Protocol-based formats also encode data in binary, but differ in that they rely on a predefined protocol definition shared between sender and receiver. Using a shared protocol frees more information out of the transmitted packet, yielding smaller payloads and faster serialization time.
@@ III-B Data Streaming Technologies
TABLE II: A comparison of different data streaming technologies.
@@ III-B3 Low Level
We do not consider other RPC systems such as Apache Flight. Instead, we rely on ProtoBuf and gRPC for their communication protocols.
A summary of the comparison of various data streaming technologies can be found in Table II. Message queue-based technologies use message queues and a publish/subscribe model to transmit data. Producers publish messages to a topic, and multiple consumers can subscribe to these topics to read messages from the queue. These systems operate in push mode, where the system delivers messages to consumers, or in pull mode, where consumers request messages from the message queuing system. RPC-based technologies define a communication protocol shared between producers and consumers, eliminating the need for an intermediate broker. Producers respond to remote procedure calls (consumer requests) to provide data. Low-level communication protocols and the ADIOS also do not require an intermediate broker. Unlike RPC technologies, they do not wait for clients’ requests to send messages, reducing communication overhead. ZeroMQ and ADIOS support zero-copy messaging transfer of raw bytes, which is particularly beneficial for large array workloads where encoding and copying data can be costly.
@@ IV Empirical Study Design
Figure 1: Illustrates the data flow from producer to consumer, indicating the places at which various performance metrics are recorded. These metrics include (A) $L_{o}$: object creation latency, (B) $T_{o}$: object creation throughput, (C) $C$: compression ratio, (D) $L_{s}$: serialization latency, (E) $L_{d}$ deserialization latency, (F) $T_{s}$: serialization throughput, (G) $T_{d}$: deserialization throughput, (H) $L_{trans}$: transmission latency, (I) $T_{trans}$: transmission throughput, (J) $L_{tot}$: total latency, and (K) $T_{tot}$: total throughput.
Figure 2: Diagram showing the architecture of our streaming framework. A Runner is used to create a Producer and Consumer pair for each type of streaming technology. Both producer and consumer are instantiated with a Marshaler that encodes data to the desired format (e.g. JSON, ProtoBuf etc.). Producers are created with a data stream object that generates data samples for transmission. Depending on the streaming method, the Consumer and Producer may connect to an external message broker.
The objective of this empirical study is to investigate and compare various streaming technologies and serialization protocols for scientific data. We examine the interplay between serialization protocol and streaming technology by exploring different combinations of them. We conduct experiments on all the technologies discussed in section III, which includes 11 different streaming technologies and 15 different serialization protocols. We test each combination of technology across eight different payloads, resulting in $11\times 15\times 8=1320$ different combinations.
@@ IV-A Performance Metrics
We consider 11 performance metrics: seven of these metrics are associated with each serialization protocol, and the remaining four are linked to the combination of streaming technology and serialization protocol. To define the metrics, we first need to establish the different sizes of data as it passes through our pipeline. We denote the size of the data straight from the stream as $S_{d}$, the size of the data after object creation as $S_{o}$, and the size of the payload after encoding to bytes as $S_{p}$. Additionally, we define the number of samples in a dataset as $N$.
To evaluate the performance of each serialization technology we measure:
@@
    Object Creation Latency ($L_{o}$) – This measures the total time taken to convert the program-specific native format (e.g. numpy array, xarray dataset) into the format required for transmission. This is an important metric because some formats, such as Capn’Proto store their data internally in a serialization-ready format. However, in reality, we often need to work with arrays that are in an analysis-ready format, such as numpy array or xarray dataset. Converting between the two models naturally incurs a penalty since it involves copying the data.
@@
    Object Creation Throughput ($T_{o}=\frac{S_{d}N}{\sum L_{o}^{(i)}}$) – This is similar to serialization and deserialization throughput, we consider object creation throughput is a sum of latencies for all samples sent) of converting a native object (e.g., numpy array, xarray dataset) to the transmission format expected by the protocol (e.g., a ProtoBuf object or a Capn’Proto object).
@@
    Compression Ratio ($C=\frac{S_{p}}{S_{o}}\times 100$) – This is defined as the ratio of the size of the payload $S_{p}$ after serialization to the size of the object $S_{o}$, for a given encoding. A smaller compression ratio ultimately means less data to be transmitted over the wire, and therefore, protocols that produce a smaller payload should be more performant.
@@
    Serialization Latency ($L_{s}$) – This is the total time taken in seconds to encode the original data into the serialized format for transmission. Encoding data with any serialization protocol incurs a non-zero cost due to the need to format, copy, and compress data for transmission. A larger serialization time can potentially negate the benefits of a smaller payload size because it increases the total transmission time.
@@
    Deserialization Latency ($L_{d}$) – This is similar to serialization time, this metric measures the total time required to deserialize a payload after transmission across the wire. As with serialization time, a slow deserialization time can also negate the effects of a smaller payload.
@@
    Serialization Throughput ($T_{s}=\frac{S_{o}}{L_{s}}$) – This is the serialization time divided by the size of the object to be transmitted. This measures how many bytes per second a serialization protocol can handle, independent of the size of the data stream.
@@
    Deserialization Throughput ($T_{d}=\frac{S_{o}N}{\sum L_{d}^{(i)}}$) – This is the deserialization time divided by the size of the object received. This measures how many bytes per second a deserialization protocol can handle, independent of the size of the data stream.
@@
    Transmission Latency ($L_{trans}$) – This is the time taken for a payload to be sent over the wire, excluding the time taken to encode the message.
@@
    Transmission Throughput ($T_{trans}=\frac{S_{d}N}{\sum L_{trans}^{(i)}}$) – This is similar to total throughput, but considers the payload size divided by the time taken to send the message over the wire, exclusive of the serialization time.
@@
    Total Latency ($L_{tot}$) – This is the total time for a payload to be transmitted from producer to consumer, inclusive of the serialization time.
@@
    Total Throughput ($T_{tot}=\frac{S_{d}N}{\sum L_{tot}^{(i)}}$) – This is the original data object size divided by the total time to send the message. Throughput measures the rate of bytes that can be communicated over the wire.
Finally, we also investigate the effect of batch size on the throughput. Grouping data into batches is a common requirement during machine learning training, and we show increasing the batch size while lowering the number of communications has a positive effect on throughput.
We make a distinction between transmission time and total time (Fig. 1). The total time is the end-to-end transmission of a message, including the time to serialize the message and send it over the wire. Transmission time is the time taken to transmit the payload excluding the serialization and deserialization times. Similarly, we can calculate total and transmission throughput.

### v2

@@ Abstract
Efficient data streaming is essential for real-time data analytics, visualization, and machine learning model training, particularly when dealing with high-volume datasets. Various streaming technologies and serialization protocols have been developed to cater to different streaming requirements, each performing differently depending on specific tasks and datasets involved. This variety poses challenges in selecting the most appropriate combination, as encountered during the implementation of streaming system for the MAST fusion device data or SKA’s radio astronomy data. To address this challenge, we conducted an empirical study on widely used data streaming technologies and serialization protocols. We also developed an extensible, open-source software framework to benchmark their efficiency across various performance metrics. Our study uncovers significant performance differences and trade-offs between these technologies, providing valuable insights that can guide the selection of optimal streaming and serialization solutions for modern data-intensive applications. Our goal is to equip the scientific community and industry professionals with the knowledge needed to enhance data streaming efficiency for improved data utilization and real-time analysis.
@@ I Introduction
With the exponential increase in data generation from large scientific experiments and the rise of data-intensive machine learning algorithms in scientific computing [1], traditional methods of data transfer are becoming inadequate. This trend necessitates efficient data streaming methods that allow end-users to access subsets of data remotely. Additionally, the drive for FAIR [2, 3], and open data mandates require that such data be publicly accessible to end users via the internet.
TABLE I: List of abbreviations used in this study.
Abbreviation Definition
ADIOS ADaptable Input Output System
AMQP Advanced Message Queuing Protocol
API Application Programming Interface
BSON Binary JSON
CBOR Concise Binary Object Representation
CCFE Culham Centre for Fusion Energy
CoAP Constrained Application Protocol
EXI Efficient XML Interchange
FAIR Findable, Accessible, Interoperable, and Reusable
gRPC Google RPC
HTTP Hypertext Transfer Protocol
JSON Javascript Object Notation
MAST Mega-Ampere Spherical Tokamak
MAST-U Mega-Ampere Spherical Tokamak Upgrade
MQTT MQ Telemetry Transport
P2P Peer-to-peer
ProtoBuf Protocol Buffers
PSON Protocol JSON
REST Representational State Transfer
RPC Remote Procedural Call
SKA Square Kilometer Array
STOMP Simple Text Orientated Messaging Protocol
TCP Transmission Control Protocol
UBJSON Universal Binary JSON
UKAEA UK Atomic Energy Authority
WAN Wide Area Network
XHTML eXtensible HyperText Markup Language
XML eXtensible Markup Language
XMPP Extensible Messaging and Presence Protocol
YAML YAML Ain’t Markup Language
ZMQ ZeroMQ
MAST [4, 5] was a fusion reactor in operation at the UKAEA, CCFE from 1999 to 2013 and its upgraded successor is MAST-U [6], which began operation in 2020. These facilities generate gigabytes of data per experimental run, with possibly many runs per day. However, the lack of public access to the archive of data produced by the MAST experiment and other fusion devices around the world has limited collaborative opportunities with international and industry partners. The pressing need to facilitate real-time data analysis [7] and leverage recent advancements in machine learning [8, 9] further emphasizes the necessity for efficient data streaming technologies. These technologies must not only handle the sheer volume of data but also integrate seamlessly with analytical tools.
In this paper, we extend the work conducted in [10] for the SKA’s radio astronomy data streaming and visualization. We explore an array of different streaming technologies available. We consider the combination of two major choices of technology when implementing a streaming service: (a) the choice of a streaming system, which performs the necessary communication between two endpoints, and (b) the choice of serialization used to convert the data into transmittable formats. Our contributions are as follows:
  • We provide a comprehensive review of 11 streaming technologies and 13 serialization methods, categorized by their underlying principles and operational frameworks.
@@
  • By testing over 143 combinations, we offer a detailed comparative analysis of their performance across eight different datasets.
@@
The rest of this work is structured as follows. Section II offers a concise review of related literature. Section III presents an overview of various serialization protocols and data streaming technologies examined in this study. Section IV outlines our experimental methodology, including the performance metrics considered, implementation specifics of our benchmark framework and the selection of datasets used for evaluation. Section V analyzes experimental results covering all performance metrics and datasets. Finally, sections VI and VII provides a discussion of the findings and recommendations of technology choices.
@@ II Related Work
Churchill et al. [7] explored using ADIOS [16] for transferring large amounts of Tokamak diagnostic data from the K-STAR facility in Korea to the NERSC and PPPL facilities in the USA for near-real-time data analysis.
We differentiate our study from these related works by evaluating 1) A wide variety of different streaming technologies, both message broker-based and RPC-based. 2) considering a large number of data serialization formats, including text, binary, and protocol-based formats. 3) We evaluate the combination of these technologies, developing an extensible framework for measuring and comparing serialization and streaming technologies. 4) Evaluating the performance over 11 different metrics. We comprehensively evaluate 11 different streaming technologies with 13 different serialization methods over 8 different datasets.
@@ III-A Serialization Protocols
TABLE II: A comparison of various serialization protocols. Type: describes how the method serializes data, e.g., text, binary, common protocol. Human readable: indicates whether the serialization scheme is legible to a human reader. Defined schema: specifies whether producer and consumer share a common knowledge of the data format prior to transmission. Code generated schema: states whether the serialization requires code to be generated from a predefined protocol.
Protocol Type Binary Human Readable Defined Schema Code Generated Schema
@@
BSON [20] Binary
UBSON [21] Binary
CBOR [22] Binary
MessagePack [23] Binary
@@
Capn’Proto (capnp, capn-packed)[27] Protocol
@@ III-A1 Text Formats
JSON [18] is another human-readable data interchange format that represents data as a collection of nested key-value pairs. JSON is commonly used for data exchange protocol in RESTful APIs. Due to the smaller payload size, it is often seen as a lower overhead alternative to XML for data interchange.
@@ III-A3 Protocol Formats
Thrift [26] is another binary data format developed by Apache Software Foundation or Apache, similar in many respects to ProtoBuf. In Thrift, data structures are also defined in a separate file, and these definitions are used to generate corresponding appropriate data structures in various supported languages. Before transmission, data are serialized into a binary format. Thrift is also designed for RPC communication and includes methods for defining services that use Thrift data structures. However, Thrift has a smaller number of supported data types compared to ProtoBuf.
@@
Additionally, we also considered the PSON format [30] and Zerializer [31]. PSON is a binary serialization format with a current implementation limited to C++ and lacks Python bindings, which restricts its applicability for our study. Zerializer, on the other hand, necessitates a specific hardware setup for implementation, placing it outside the scope of our study due to practical constraints. Consequently, while these formats might offer potential advantages, their limitations in terms of language support and hardware requirements precluded their inclusion in our experimental evaluation.
A summary of serialization protocols can be found in Table II. The text-based formats represent data using a text-based markup. While human-readable, text-based formats suffer from larger payload sizes and serialization costs due to the overhead of the markup describing the data. In contrast, binary formats serialize the data to bytes before transmission. These formats are not human-readable, but achieve a better payload size with lower serialization costs. Protocol-based formats also encode data in binary, but differ in that they rely on a predefined protocol definition shared between sender and receiver. Using a shared protocol frees more information out of the transmitted packet, yielding smaller payloads and faster serialization time.
@@ III-B Data Streaming Technologies
TABLE III: A comparison of different data streaming technologies.
@@ III-B3 Low Level
We do not consider other RPC systems such as Apache Arrow Flight [43] which rely on ProtoBuf and gRPC for their underlying communication protocols.
A summary of the comparison of various data streaming technologies can be found in Table III. Message queue-based technologies use message queues and a publish/subscribe model to transmit data. Producers publish messages to a topic, and multiple consumers can subscribe to these topics to read messages from the queue. These systems operate in push mode, where the system delivers messages to consumers, or in pull mode, where consumers request messages from the message queuing system. RPC-based technologies define a communication protocol shared between producers and consumers, eliminating the need for an intermediate broker. Producers respond to remote procedure calls (consumer requests) to provide data. Low-level communication protocols and the ADIOS also do not require an intermediate broker. Unlike RPC technologies, they do not wait for clients’ requests to send messages, reducing communication overhead. ZeroMQ and ADIOS support zero-copy messaging transfer of raw bytes, which is particularly beneficial for large array workloads where serializing and copying data can be costly.
@@ IV Empirical Study Design
Figure 1: Illustrates the data flow from producer to consumer, indicating the places at which various performance metrics are recorded. These metrics include (1) $L_{o}$: object creation latency, (2) $T_{o}$: object creation throughput, (3) $C$: compression ratio, (4) $L_{s}$: serialization latency, (5) $L_{d}$ deserialization latency, (6) $T_{s}$: serialization throughput, (7) $T_{d}$: deserialization throughput, (8) $L_{trans}$: transmission latency, (9) $T_{trans}$: transmission throughput, (10) $L_{tot}$: total latency, and (11) $T_{tot}$: total throughput.
Figure 2: Diagram showing the architecture of our streaming framework. A Runner is used to create a Producer and Consumer pair for each type of streaming technology. Both producer and consumer are instantiated with a Marshaler that encodes data to the desired format (e.g. JSON, ProtoBuf, etc.). Producers are created with a data stream object that generates data samples for transmission. Depending on the streaming method, the Consumer and Producer may connect to an external message broker.
The objective of this empirical study is to investigate and compare various streaming technologies and serialization protocols for scientific data. We examine the interplay between serialization protocol and streaming technology by exploring different combinations of them. We conduct experiments on all the technologies discussed in section III, which includes 11 different streaming technologies and 13 different serialization protocols. We test each combination of technology across eight different payloads (data types), resulting in $11\times 13\times 8=1144$ different combinations.
Finally, in we also investigate the effect of batch size on the throughput. Grouping data into batches is a common requirement during machine learning training, and we show increasing the batch size while lowering the number of communications has a positive effect on throughput.
@@ IV-A Performance Metrics
We evaluate 11 performance metrics. The first seven metrics are associated with serialization protocol, and the remaining four are linked to the combination of streaming technology and serialization protocol. To define these metrics, we first establish the different data sizes at various stages of the pipeline. $S_{d}$ denote the size of the original data, $S_{o}$ the size after serialization (e.g., the size after creating a gRPC object), and $S_{p}$ the size of the payload after serializing it to bytes. Additionally, we define the number of samples in the dataset as $N$.
The metrics we consider are:
@@
    Object Creation Latency ($L_{o}$) measures the total time taken to convert the program-specific native format (e.g. Numpy array, Xarray dataset) into the format required for transmission. This is an important metric because some formats, such as Capn’Proto store their data internally in a serialization-ready format. However, in reality, we often need to work with arrays that are in an analysis-ready format, such as Numpy array or Xarray dataset. Converting between the two models can incur a penalty since it may involve copying the data.
@@
    Object Creation Throughput ($T_{o}=\frac{S_{d}}{L_{o}}$) measures the total size of a native data type $S_{d}$ (e.g., Numpy array, or Xarray dataset) divided by the total time to convert it to the transmission format expected by the protocol (e.g., a ProtoBuf object or a Capn’Proto object).
@@
    Compression Ratio ($C=\frac{S_{p}}{S_{o}}\times 100$) is defined as the ratio of the size of the payload $S_{p}$ after serialization to the size of the object $S_{o}$. A smaller compression ratio ultimately means less data to be transmitted over the wire, and therefore, protocols that produce a smaller payload should be more performant.
@@
    Serialization Latency ($L_{s}$) is the total time taken to encode the original data into the serialized format for transmission. Serializing data with any protocol incurs a non-zero cost due to the need to format, copy, and compress data for transmission. A larger serialization time can potentially negate the benefits of a smaller payload size because it increases the total transmission time.
@@
    Deserialization Latency ($L_{d}$) is similar to serialization time, this metric measures the total time required to deserialize a payload after transmission across the wire. As with serialization time, a slow deserialization time may negate the effects of a smaller payload.
@@
    Serialization Throughput ($T_{s}=\frac{S_{o}}{L_{s}}$) is the serialization time divided by the size of the object to be transmitted. This measures how many bytes per second a serialization protocol can handle, independent of the size of the data stream.
@@
    Deserialization Throughput ($T_{d}=\frac{S_{o}}{L_{d}}$) is the deserialization time divided by the size of the object received. This measures how many bytes per second a deserialization protocol can handle, independent of the size of the data stream.
@@
    Transmission Latency ($L_{trans}$) is the time taken for a payload to be sent over the wire, excluding the time taken to encode the message.
@@
    Transmission Throughput ($T_{trans}=\frac{S_{d}}{L_{trans}}$) is similar to total throughput, but considers the payload size divided by the time taken to send the message over the wire, exclusive of the serialization time.
@@
    Total Latency ($L_{tot}$) is the total time for a payload to be transmitted from producer to consumer, inclusive of the serialization time.
@@
    Total Throughput ($T_{tot}=\frac{S_{d}}{L_{tot}}$) is the original data object size divided by the total time to send the message. Throughput measures the rate of bytes that can be communicated over the wire.
We make a distinction between transmission time and total time (Figure 1). The total time is the end-to-end transmission of a message, including the time to serialize the message and send it over the wire. Transmission time is the time taken to transmit the payload excluding the serialization and deserialization times. Similarly, we can calculate total and transmission throughput.